OpenStudy (anonymous):

f^-1(4) when f(x) = 3x minus 4, all over 5 will give medals, please help me solve this

OpenStudy (ipwnbunnies):

f^(-1)(x) is the inverse of f(x). To find the inverse: 1. Replace f(x) with y 2. Swap the places of x and y 3. Solve for y again 4. Replace y with f^(-1)(x) 5. Plug in your x value

OpenStudy (anonymous):

y= 3x - 4 over 5 x= 3y - 4 over 5 4= 3y - 4 over 5 am I right so far @iPwnBunnies

OpenStudy (ipwnbunnies):

That's another way to do it lol. I said solve for y before you plug in the x-value to distinguish f(x) from f^(-1).

OpenStudy (anonymous):

Then times five by four and add 4 so 24= 3y then divide by 3 and I get 8?

OpenStudy (anonymous):

Oh I'm sorry I get a bit confused

OpenStudy (ipwnbunnies):

It is 8 though, when I do it quick in my head.

OpenStudy (ipwnbunnies):

x = (3y-4)/5 y = (5x+4)/3 So, f^(-1)(x) = (5x+4)/3 Those were the extra steps to distinguish the two functions lol.

OpenStudy (anonymous):

Ok thank you so much! (: